OpenStudy (lgbasallote):

find the volume generated by rotating the region bounded by \(y = x^2 +1\), y = 5 and x = 0 about y = -1

OpenStudy (lgbasallote):

i got \[\frac{112\pi}{15}\]

OpenStudy (lgbasallote):

is that right @dpaInc ?

OpenStudy (anonymous):

i got 704pi/15...

OpenStudy (lgbasallote):

you used dx right?

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

washer...

OpenStudy (anonymous):

outer radius is 6 inner radius is x^2 + 1 + 1 = x^2 +2

OpenStudy (lgbasallote):

washer? o.O

OpenStudy (lgbasallote):

isnt it cylindrical shell??

OpenStudy (anonymous):

if you're going dx, then your doing washer. if you're going dy, then you can use shell. remember the axis of rotation is y=-1 (horizontal)

OpenStudy (lgbasallote):

according to my teacher